OpenStudy (anonymous):

Negative exponents question, don't get a simple question because teacher taught it wrong. 3 to the negative power of 5?

jimthompson5910 (jim_thompson5910):

anything to a negative power can be written as the reciprocal of the base to the positive exponent. So \[3^{-5}=(\frac{3}{1})^{-5}=(\frac{1}{3})^5=\frac{1^5}{3^5}=\frac{1}{243}\]